Solving quadratic equations by factoring. N square + 7n+15=5
Answer by ankor@dixie-net.com(22740)

Solving quadratic equations by factoring. N square + 7n+15=5
n^2 + 7n + 15 = 5
n^2 + 7n + 15 - 5 = 0
n^2 + 7n + 10 = 0
Factors to
(n+5)(n+2) = 0
two solutions
n = -2
n = -5